OpenStudy (jkbo):

Please help What is the solution of matrix equation [2 5] [3] [1 3] x= [-1]

OpenStudy (anonymous):

does that x have brackets around it?

OpenStudy (jkbo):

No it doesnt

OpenStudy (anonymous):

ok... hmm. That should be a matrix with two rows and one column there. So I would set it up like\[x=\left[\begin{matrix}x_1 \\ x_2 \end{matrix}\right]\] and then solve just like the other one we did. You're answer will be a matrix with 2 rows and one column

OpenStudy (jkbo):

alright and I got [14] [ -5]

OpenStudy (anonymous):

Great job!

OpenStudy (jkbo):

I was wondering if thats correct

OpenStudy (anonymous):

it absolutely is!

OpenStudy (jkbo):

Thank you

OpenStudy (jkbo):

@caseyrt can you help me with this question ? Which matrix equation has the solution X = [3] [-5]

OpenStudy (anonymous):

sure, are there some answer options?

OpenStudy (jkbo):

A.[9 5] [2] [7 -4] x =[1] B. [-9 5] [2] [7 4] x =[1] C.[9 5] [2] [7 4] x =[1] D. [9 5] [2] [-7 4] x =[1]

OpenStudy (anonymous):

ok so we just need to put the little matrix they give us in the place of x in each place and see which one comes out being true! ...

OpenStudy (anonymous):

so for the first one, 9*3-5*5=2

OpenStudy (anonymous):

so the top is correct

OpenStudy (anonymous):

but the bottom is 7*3+4*5=41 which is obviously not 1 like it says it should be, so A isn't it

OpenStudy (anonymous):

but the top part of A was right, so the answer will have its top part the same as A's, which means the answer isn't B, either. Of the two that are left, can you find out which one has the correct bottom row?

OpenStudy (jkbo):

it would be .d

OpenStudy (anonymous):

d would give us -7*3-4*5... which I think is not 1... do you see why that is?

OpenStudy (jkbo):

Yes because that equals -41

OpenStudy (anonymous):

yah I guess I meant to ask do you see why I set up the numbers the way I did?

OpenStudy (anonymous):

because that's really the only hard thing about these problems... but it can be pretty tough!

OpenStudy (jkbo):

Alright so the answer would be C. then

OpenStudy (anonymous):

right.

OpenStudy (jkbo):

Thanks